Patterns Complete the pattern 2,3,5,9,?,?,?,
Answer by mathie123(224)

We have the pattern
2,3,5,9,?,?,?,
If we look at what we have added each time we add 1 then 2 then 4... no real pattern there.
After some trying out patterns we see that
2*2-1=3
3*2-1=5
5*2-1=9

So we have a pattern -- the previous number times 2 minus 1.

So the next term is 9*2-1=17
and the term after is
17*2-1=33


I will let you figure out the next term


Hopefully this helps!